OpenStudy (anonymous):

Graphing help, cant find where to put my points. Solve the system of equations by graphing.

OpenStudy (anonymous):

x + y = 0 5x − 2y = 14 (x,y) x=2 y=-2.0 cannot figure out my graph points

OpenStudy (anonymous):

the first two r the lines, what is the third one?

OpenStudy (anonymous):

From the graph, it appears that the intersection is at the point: (x, y) = (2, -2.0)

OpenStudy (anonymous):

so u want to draw the lines?

OpenStudy (anonymous):

yes sorry i just saw the origianl question did not post.. Plot the equations x + y = 0 and 5x − 2y = 14 below.

OpenStudy (anonymous):

the first one is x= -y line which passes through (0,0) n dividing 2nd and 3rd quadrant

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x%2By%3D0

OpenStudy (anonymous):

5x-2y=14 x intercept=14/5 y intercept is -7 joining both we get the line

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=5x-2y%3D14

OpenStudy (anonymous):

thanks for the website...

OpenStudy (anonymous):

when u draw the line in same graph, u get the point of intersection